OpenStudy (anonymous):

What is the y-intercept of the equation y=0.5(4)x?

jimthompson5910 (jim_thompson5910):

I'm assuming that this is \[\Large y = 0.5(4)^{x}\] correct?

OpenStudy (anonymous):

yes i think so

jimthompson5910 (jim_thompson5910):

is that x above the 4 a bit?

jimthompson5910 (jim_thompson5910):

If so, then... Plug in x = 0 and evaluate to find y \[\Large y = 0.5(4)^{x}\] \[\Large y = 0.5(4)^{0}\] \[\Large y = 0.5(1)\] \[\Large y = 0.5\] So the y-intercept is (0, 0.5)

OpenStudy (anonymous):

is it zero

OpenStudy (anonymous):

plane curve | Cartesian equation y = 0.5 4 x | y-intercept

OpenStudy (anonymous):

i have no clue how to do this and im i the middle of a quiz

OpenStudy (anonymous):

properties of line x-intercept | 0 y-intercept | 0 slope | 2.

OpenStudy (anonymous):

Wolframalpha is a great resource, but it's best to lead the student to the solution via a more step-by-step application approach. You quote wolframalpha directly in every single answer.

OpenStudy (anonymous):

yes that link i had already sented

OpenStudy (anonymous):

ya so the answers is 0

OpenStudy (anonymous):

jim_thompson5910 has the correct approach. See what is going on there.

OpenStudy (anonymous):

ok so the answer is 0.5

OpenStudy (anonymous):

Yes, the y-intercept is at y = 0.5

OpenStudy (anonymous):

ok thank you guys